A quadratic polynomial of the form y=x^2+bx+c has x-intercepts at x=-7 and x=-3. What is the y-intercept of this parabola?
If anyone can help me solve this, I will gladly appreciate it.

Answer by ikleyn(52781)

.
A quadratic polynomial of the form y=x^2+bx+c has x-intercepts at x=-7 and x=-3.
What is the y-intercept of this parabola?
If anyone can help me solve this, I will gladly appreciate it.
~~~~~~~~~~~~~~~~~~~~

y-intercept is the value of this quadratic polynomial at x= 0.


Substitute x= 0 into the polynomial, and you will see that y-intercept is the value
of the coefficient "c".


According to Vieta's theorem, coefficient "c" (the constant term) is the product of the roots of the polynomial.


The roots are -7 and -3 (given), and their product is (-7)*(-3) = 21.


So, y-intercept is 21.    ANSWER

Solved.